OpenStudy (anonymous):

Solve. a-5/a+4=5/4 a = 0 a = -1 a = -9 a = -40

OpenStudy (kendricklamar2014):

Step 1. Cross Multiply: (a−5)*(4)=5*(a+4) 4a−20=5a+20

OpenStudy (anonymous):

thank you

OpenStudy (kendricklamar2014):

Step 2. Subtract 5a from both sides. 4a−20−5a=5a+20−5a −a−20=20

OpenStudy (kendricklamar2014):

Step 3: Add 20 to both sides. −a−20+20=20+20 −a=40

OpenStudy (kendricklamar2014):

Step 4: Divide both sides by -1 -a/-1 = 40/-1 a=−40

OpenStudy (kendricklamar2014):

yw :)

OpenStudy (anonymous):

thanks:)

OpenStudy (kendricklamar2014):

=)